OpenStudy (anonymous):

(3sqrt7-9sqrt3)(4sqrt7+12sqrt3)

OpenStudy (anonymous):

\[(3\sqrt{7}-9\sqrt{3})(4\sqrt{7}+12\sqrt{3})\]

OpenStudy (anonymous):

\[12\sqrt{7}+36\sqrt{21}-36\sqrt{21}\]

OpenStudy (anonymous):

\[12\sqrt{7}\]

OpenStudy (anonymous):

hey \[\sqrt{7} x \sqrt{7} =7\]

OpenStudy (anonymous):

you sure about this ans??

OpenStudy (anonymous):

i dont know thats why i am asking lol

OpenStudy (anonymous):

\[(x-7)^{2}=20 is this answer correct 7+2\sqrt{5},7-2\sqrt{5}\]